Wesley invested $6500 in a mutual fun at a yearly rate of 3 3/4%. If he has earned $975 in interest, how long has the money been invested?
A. 2.5 years B. 3 years C. 4 years D. 4.5 years
: Wesley invested $6500 in a mutual fun at a yearly rate of 3 3/4%. If he has earned $975 in interest, how long has the money been invested?
A. 2.5 years B. 3 years C. 4 years D. 4.5 years

Answer by mangopeeler07(441)

3.75% of 6500

x/6500=3.75/100

Cross multiply
100x=24375

Divide by 100
x=243.75

243.75 each year for x years=$975

975/243.75=x
4=x

Answer:
C) 4 years